OpenStudy (anonymous):

If the sin 30° is 1 over 2, then the cos ____° = _____.

OpenStudy (anonymous):

60° ; 1 over 2 30° ; square root 2 over 2 60° ; square root 3 over 2 30° ; 1

OpenStudy (anonymous):

My guess is C.

OpenStudy (anonymous):

I know sin 30° 1/2 -> cos ?° ?/2

OpenStudy (ipwnbunnies):

\[\sin(\theta) = \cos(90 - \theta)\]

OpenStudy (anonymous):

So either A or C?

OpenStudy (anonymous):

A? 30:60 1:2 ? Am I right?

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (ipwnbunnies):

Yes. The sin30 degrees is 1/2. So is the cos60 degrees.

OpenStudy (anonymous):

Thanks now I understand sin to cos.